OpenStudy (anonymous):

Can someone explain how to take a derivate when there is a sqrt. The sqrt throws me off and I have a test coming up and want to really understand this. 1.) xsqrt(x+3)

OpenStudy (anonymous):

Have you tried writing it as a fractional exponent?

OpenStudy (anonymous):

That is to say \[\sqrt{x+3} = (x+3)^\frac{1}{2}\]

OpenStudy (anonymous):

so then I would just multiply it by the 1/2

OpenStudy (anonymous):

Exactly! Of course, you'll want to make sure that you also multiply the expression by the derivative of what's inside the parentheses.

OpenStudy (anonymous):

Okay that makes much more sense! thanks so much!

OpenStudy (anonymous):

You bet. What did you get for your final answer?